OpenStudy (anonymous):

Let f(x)= 5, g(x)=-3 x - 1, and h(x)=-4 x^2. Consider the inner product \displaystyle = integral from 0 to 4 p(x)q(x) dx in the vector space C^0[0,4]. Use the Gram-Schmidt process to determine an orthonormal basis for the subspace of C^0[0,4] spanned by the functions f(x), g(x), and h(x).... help??

OpenStudy (anonymous):

opps sorry copied the \displaystyle. any ideas of where to start?

OpenStudy (anonymous):

have done gram-schmidt before with vectors and only remember that it is a real pain. don't know how one does it with inner product given as integral. sorry

OpenStudy (anonymous):

let me look in taylor

OpenStudy (anonymous):

using functions makes things a little weird thats what throws me for a loop

OpenStudy (anonymous):

so are we thinking of f, g and h as vectors? then first step is to normalize them yes?

OpenStudy (anonymous):

\[\int_0^2 f^2dx=\int_0^4 25dx=100\]

OpenStudy (anonymous):

so normalized we get \[v_1(x)=\frac{1}{25}\] hope this is the right track. maybe i am all wrong

OpenStudy (anonymous):

so your normalizing by squaring the function and dividing by it. so far seems to be equivalent to a unit vector

OpenStudy (anonymous):

so what would my vectors be for f(x) g(x) h(x)? i guess if I have these I could solve similarly to the normal gram-schmidt way.

OpenStudy (anonymous):

oh i see f(x) would be (1 0 0) g(x) is (-1 -3 0) ect ect

OpenStudy (anonymous):

rather (5 0 0)

OpenStudy (anonymous):

ok this is what i have written in taylor: If A_1, A_2, A_3 we start by defining \[v_1=\frac{A_1}{|A_1|}\] and let\[ B_2 = A_2 -\text{ the projection of} A_1 \text{on } v_1\]

OpenStudy (anonymous):

in our case i guess \[A_1=5\] so \[v_1=\frac{1}{25}\]

OpenStudy (anonymous):

now \[B_2=A_2-(A_2\circ v_1)v_1\] where circle means the integral.

OpenStudy (anonymous):

so it looks like \[B_2=g(x)-(\int_0^4g(x)\times5)\times \frac{1}{25}\]

OpenStudy (anonymous):

if i am reading this thing correctly.

OpenStudy (anonymous):

inner product means integral and i am just writing down what the process is. i could be way way off

OpenStudy (anonymous):

well lets just check

OpenStudy (anonymous):

what would our first vector basis be?

OpenStudy (anonymous):

I think it's off. . .

OpenStudy (anonymous):

\[\frac{1}{25}\] which is \[f(x)=5\] normalized

OpenStudy (anonymous):

we should get a (v1, v2, v3) such that these are the vectors spanning our basis

OpenStudy (anonymous):

1/25 is incorrect

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you need more than one answer for gram schmidt

OpenStudy (anonymous):

no worries i appreciate the help so far

OpenStudy (anonymous):

yes we understand that, but we should get an answer for each step, thus 3 vectors spanning our basis

OpenStudy (anonymous):

no just one step at a time. \[\int_0^4 5^2dx = 100\] yes? i thought the first vector in your normalization process was just the first vector normalized. then you go from there. do you happen to know what the first vector is?

OpenStudy (anonymous):

first vector is just a constant and norm is the integral so when you divide you get a vector of length one yes?

OpenStudy (anonymous):

in normal gram schmidt you normal the first vector, this is your first unit vector, which is the first of the (v1, v2, ...vn) vectors that are orthonormal

OpenStudy (anonymous):

Sorry to but in. . .I have my lin alg book next to me now.

OpenStudy (anonymous):

by all means how does it work with functions because i cant figure it out

OpenStudy (anonymous):

please help because i have only done this with vectors written as coordinates

OpenStudy (anonymous):

Ok. . .

OpenStudy (anonymous):

oh i see a mistake on my part. yikes. integral is 100 but norm is square root of that, so 10

OpenStudy (anonymous):

I think you were doing the right thing. . . <u1, v1> is the def integral of the first 2 functions multiplied together

OpenStudy (anonymous):

so normalized first vector is \[v_1=\frac{1}{2}\]

OpenStudy (anonymous):

thats correct

OpenStudy (anonymous):

now let me try to find \[B_2\] which is \[B_2=A_2-(A_2\circ v_1)\times v_1\]

OpenStudy (anonymous):

\[B_2=g(x)-(\int_0^4g(x)\times \frac{1}{2})dx\times {1}{2}\]

OpenStudy (anonymous):

should be times 1/2 not times 12

OpenStudy (anonymous):

by the way, where are you getting this from so I can understand the process

OpenStudy (anonymous):

i am looking at a copy of Taylor advanced calc

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

the integral is -7 so i get \[B_2=-3x-8\]

OpenStudy (anonymous):

now we need \[\frac{B_2}{||B_2||}\] and \[||B_2||=\sqrt{\int_0^4 (-3x-8)^2dx}\]

OpenStudy (anonymous):

i get 832 for the integral and \[8\sqrt{13}\] for the norm

OpenStudy (anonymous):

mathgirls still there? did i mess up yet?

OpenStudy (anonymous):

shes gone

OpenStudy (anonymous):

which means we get -3x-8/8sqrt(13)

OpenStudy (anonymous):

that is what i get, but really i would not bet my beer money on it

OpenStudy (anonymous):

do you have any idea what the answer is?

OpenStudy (anonymous):

you seemed to know that the first was 1/2

OpenStudy (anonymous):

its a webwork with unlimited attempts so all i can do is try, but i have no idea what the sanwer is nor can i find a decent example to explain the process

OpenStudy (anonymous):

can you try that answer? because there is no use in continuing if that is wrong

OpenStudy (anonymous):

its not correct

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ill figure it out. no worries i appreciate the help!

OpenStudy (anonymous):

sorry. it is \[g(x)=-3x-1\] yes?

OpenStudy (anonymous):

yeah i tried 8sqrt(13)*(-3x-1)

OpenStudy (anonymous):

integral is right i checked it

OpenStudy (anonymous):

integral is -14 then multiply by \[v_1\] which gives is 1/2 giving -7 and oh damn!

OpenStudy (anonymous):

it says SUBTRACT what a bush league error.

OpenStudy (anonymous):

lordamercy it is \[g(x)-(-7)=g(x)+7=-3x+6\]

OpenStudy (anonymous):

ok now lets try it with that. the norm is \[\sqrt{\int_0^4(-3x+6)^2dx}\]

OpenStudy (anonymous):

integral is 48 norm is \[4\sqrt{3}\]

OpenStudy (anonymous):

making \[v_2=\frac{-3x+6}{4\sqrt{3}}\]

OpenStudy (anonymous):

yes nice

OpenStudy (anonymous):

what an elementary mistake. subtracting -7 by subtracting 7. try that and see if we hae better luck

OpenStudy (anonymous):

yeah its correct

OpenStudy (anonymous):

whew. now i feel like somewhat less of an idiot. ok one more to go yes?

OpenStudy (anonymous):

yep. ill definitely have to review this process, it is not like the regular gram- schmidt

OpenStudy (anonymous):

\[B_3=A_3-(A_3\circ v_1)v_1-(A_3\circ v_2)v_2\]

OpenStudy (anonymous):

just so i am understanding the terminology here, b3 is my projection of my third vector onto my second?

OpenStudy (anonymous):

\[A_3=h(x)=-4x^2\] \[(A_3\circ v_1)=\int_0^4(-4x^2)(\frac{1}{2})dx\times \frac{1}{2}\]

OpenStudy (anonymous):

\[B_3\] is orthogonal to \[v_1\] and \[v_2\]

OpenStudy (anonymous):

finally \[(A_3\circ v_2)v_2=\int(-4x^2)(\frac{-3x+6}{4\sqrt{3}})dx\times (-3x+6)\]

OpenStudy (anonymous):

\[(A_3 \circ v_1)=\int_0^4x^2dx=\frac{64}{3}\]

OpenStudy (anonymous):

this is a pain just like i remembered

OpenStudy (anonymous):

well if i could give you 50 medals i definitely would lol

OpenStudy (anonymous):

\[\int_0^4 \frac{3x^3-6x^2}{\sqrt{3}}dx=\frac{64\sqrt{3}}{3}\]\]

OpenStudy (anonymous):

multiply by \[-3x+6\] so get \[64\sqrt{3}x+128\]

OpenStudy (anonymous):

so \[B_3=-4x^2-\frac{64}{3}-64\sqrt{3}x-128\]

OpenStudy (anonymous):

guess we could clean that up.

OpenStudy (anonymous):

i get \[-4x^2-64\sqrt{3}x-\frac{448}{3}\] but by now i could have made a mistake.

OpenStudy (anonymous):

and finally... we need \[\sqrt{\int_0^4 (-4x^2-64\sqrt{3}x-\frac{448}{3})^2dx}\]

OpenStudy (anonymous):

i get something really ugly.

OpenStudy (anonymous):

lemme try

OpenStudy (anonymous):

oh damn damn damn

OpenStudy (anonymous):

uh oh a mistake somewhere?

OpenStudy (anonymous):

oh no i thought i caught another stupid error but it is ok

OpenStudy (anonymous):

i wolframed it and got approx 837

OpenStudy (anonymous):

hold on i may have made another stupid error

OpenStudy (anonymous):

another goddamn minus sign

OpenStudy (anonymous):

\[(A_3\circ v_2)=-64\sqrt{3}x+128\]

OpenStudy (anonymous):

so when we SUBTRACT we get \[B_3=-4x^2+64\sqrt{3}x-128-\frac{64}{3}\]

OpenStudy (anonymous):

\[B_3=-4x^2+64\sqrt{3}x-\frac{448}{3}\]